OpenStudy (anonymous):

Rectangle A has an area of 16 − x^2. Rectangle B has an area of x^2 + 2x − 24. In simplest form, what is the ratio of the area of Rectangle A to the area of Rectangle B ? Show your work. What is the value, in centimeters, of x?

OpenStudy (anonymous):

factorize 16-x^2 = (4+x) (4-x) similarly the other eqn x^2+2x-24 = (x-4) (x+6) now take the ratio